If a ½ hp pump runs for 21 minutes, what is the energy used in kW-hr?


1
Expert's answer
2021-10-13T09:08:07-0400

1 HP = 0.75 kW

"\frac{0.75}{2} = 0.375 \;kW \; per \; 21\; min"

Proportion:

0.375 kW – 21 min

x kW – 60 min

"x = \frac{0.375 \times 60}{21} = 1.071 \;kW-hr"

Answer: 1.071 kW-hr